# PPFPL: Cross-silo Privacy-preserving Federated Prototype Learning Against Data Poisoning Attacks on Non-IID Data

### v4

@@ VII Experiments
VII-A Experimental Settings
VII-A1 Datasets and Models
Similar to previous works [16][36][52], we utilize three public available datasets, namely MNIST, FMNIST, and CIFAR10, to evaluate performance of our PPFPL.
  1. 1.
    MNIST: This dataset comprises handwritten digits, with a training set of 60,000 samples in 10 classes and a test set of 10,000 samples, each represented as a 28 $\times$ 28 gray-scale image.
  2. 2.
    FMNIST: The FMNIST is a clothing image dataset containing a training set of 60,000 samples in 10 classes and a test set of 10,000 samples, each of which is a 28$\times$28 grey-scale image.
  3. 3.
    CIFAR10: As a color image dataset, CIFAR10 contains a training set of 50,000 samples in 10 classes and a test set of 10,000 samples, with each image measuring 32$\times$32 pixels.
Furthermore, we apply Convolution Neural Network (CNN) as local model to both MNIST and FMNIST. For CIFAR-10, we employ ResNet18 as the local model, initialized with pre-trained parameters. These initial parameters have an accuracy of 27.5% on CIFAR10’s test set.
VII-A2 Hyper-parameters Settings of FL
We employ a cross-silo configuration in FL setup. Specifically, we set up 20 clients, each of which uploads local prototypes at each communication round. The number of communication rounds is set to 100, 150, and 150 for MNIST, FMNIST, and CIFAR10, respectively. We configure the local learning rate to $\eta=0.01$, the importance weight to $\lambda=1$, a batch size to 64, and the number of local iterations to $E=5$. Except for special declared, the above default settings are applied. These hyper-parameters are consistent across all clients.
VII-A3 Non-IID Settings
To simulate the Non-IID data in cross-silo FL, we create class-space heterogeneity among clients, which is common in the cross-silo scenarios. Specifically, large organizations (e.g., hospitals, companies) possess different data classes, and their class distributions may differ significantly, or even be missing some classes altogether. When these organizations participate in federated training, the union of their data classes defines the entire FL classification task. This phenomenon leads to Non-IID data across organizations. The Dirichlet distribution assumes that each client’s data is sampled from all classes, meaning each client typically have a positive percentage of every class. However, in cross-silo scenarios, some clients may have no samples from certain classes, which contradicts the assumptions of Dirichlet distribution.
To model the data distribution in cross-silo scenarios, similar to previous works [22][53], we define Avg as the mean number of data classes per client, and Std as the standard deviation of these class counts. In our experiments, we fix Avg to be 3 , 4 or 5, and fix Std to be 1 or 2, aiming to create the class-space heterogeneity. Clients are randomly assigned classes, with partial class overlap among them. To visualize the different data distributions, we plot heat maps as shown in Fig. 2.
@@ VII-A3 Non-IID Settings
Figure 2: Heat maps for different data distributions under heterogeneity distribution.
VII-A4 Setting of Data Poisoning Attacks
@@
In our experiments, we consider two types of data poisoning attacks: feature attacks and label attacks. For feature attacks, malicious clients randomly alter their own training data features in a completely randomized manner without following any specific rule. For label attacks, malicious clients alter the labels of their own training data to incorrect labels. All training data of malicious clients are modified. Malicious clients do not dynamically change their behaviors during federated training, and the attack remains consistent throughout the training period. Since the number of malicious clients affects FL performance differently, we set different proportions of malicious clients. Formally, Att is set as the proportion of malicious clients.
VII-A5 Evaluation Measure
@@
Our goal is to improve the performance of cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks. Therefore, we evaluate the FL performance by testing the average accuracy of benign clients. For experiment comparison, we use FedAvg as the baseline, assuming that FedAvg has not suffered data poisoning attacks. In addition, we compare our PPFPL with robust schemes (i.e., Krum [35], Foolsgold [54], and FedDMC [36]) and privacy-preserving robust schemes (i.e., ShieldFL [18], PBFL [16], and AntidoteFL [20]) under the same experiment conditions.
VII-B Experimental Results
VII-B1 Visualization of Prototypes
@@
We visualize the prototypes submitted by clients in PPFPL using t-SNE, as shown in Fig. 3. Specifically, we display the distribution of prototypes from clients under feature and label attacks with $\textit{Att}=20\%$ on the CIFAR10 dataset within a single communication round. We observe that the prototypes of malicious clients deviate significantly from those of benign clients, regardless of the features attacks or labels attacks. In addition, we notice that the prototypes submitted by malicious clients are concentrated in a ring area. This is because when malicious clients perform local model training, the features or labels are disrupted, causing the directions of the generated prototypes to spread from the center point to the surrounding areas. Most importantly, we observe a clear difference between the prototypes submitted by benign and malicious clients. This difference is prominent because it is independent of the data distribution. The finding validates the effectiveness of using prototypes to defend against data poisoning attacks. To filter out these malicious prototypes, it is essential to employ a secure aggregation protocol that ensures Byzantine-robust aggregation results. The following is the security evaluation of PPFPL.
((a)) Feature attack,$\textit{Avg}=3,\textit{Std}=2$.
((b)) Feature attack,$\textit{Avg}=5,\textit{Std}=1$.
((c)) Label attacks, $\textit{Avg}=3$, $\textit{Std}=2$.
((d)) Label attacks, $\textit{Avg}=5$, $\textit{Std}=1$.
@@
Figure 3: The t-SNE visualization of prototypes submitted by benign and malicious clients: different colors indicate distinct classes, where $\cdot$” denotes prototypes from benign clients, and “$\times$” represents those from malicious clients.
VII-B2 Security Evaluation
@@
To evaluate the security of our framework, we test PPFPL’s performance against feature attacks and label attacks on three datasets under Non-IID setting with $\textit{Avg}=3$ and $\textit{Std}=2$. Specifically, the proportion of malicious clients is set to Att = 20% and the detection threshold of PPFPL is configured as $\chi=0$. Furthermore, we compare PPFPL with existing schemes (i.e., FedAvg, ShieldFL, PBFL, FedDMC, and AntidoteFL), where the global model learned by FedAvg is not subject to data poisoning attacks.
The experiment results are shown in Fig. 4, where the black dashed line represents the highest accuracy of PPFPL without attacks. We observe that the accuracy of ShieldFL and PBFL (under attacks) is severely lower than FedAvg (no attack) in CIFAR10. This indicates that as the dataset becomes more complex, the negative impact of data poisoning attacks on model accuracy becomes more obvious. This is mainly because the complex dataset increases the difficulty of defense, and poisoned Non-IID data degrades the defense performance. Notably, FedDMC and AntidoteFL achieve higher performance compared to FedAvg because they adjust the aggregation weights of the model updates, which mitigates the bias caused by Non-IID data and improves the performance of federated learning. Moreover, PPFPL (under attack) outperforms FedAvg (no attacks), with its accuracy approaching the highest accuracy of PPFPL without attacks when it nears convergence. This demonstrates that PPFPL guarantees the high accuracy of the learned model in the presence of poisoned Non-IID data, which is benefited that the prototype is not affected by tampered data distribution, while the secure aggregation protocol resists malicious prototypes submitted by clients. Thus, our PPFPL overcomes the difficulty confronted by these defense schemes, and improves the FL performance in poisoned Non-IID data.
Additionally, we evaluate the change of loss during training for PPFPL and the compared schemes, as shown in Fig. 5. We notice that the loss of PBFL fluctuates significantly during training process, while PPFPL is relatively smooth. This indicates that data poisoning attacks disrupt the convergence of PBFL, while the convergence of PPFPL is not affected by data poisoning attacks, thus ensuring the accuracy of the model. This is consistent with our Corollary 3. Specifically, PPFPL still satisfies convergence in the presence of data poisoning attacks as long as $\lambda$, $E$, and $\eta$ satisfy a specific relationship among them. Moreover, PPFPL’s loss at convergence is lower than other defense schemes under CIFAR10, which can satisfy the design goal of security. Notably, our experiments show that the performance impact of feature attacks and label attacks is similar for PPFPL and other schemes. This indicates that the impact on model performance is similar under poisoned Non-IID data, regardless of the type of data poisoning attacks.
((a)) Feature attacks
((b)) Label attacks
Figure 6: Stability of PPFPL under CIFAR10 against data poisoning attacks.
VII-B3 Different Data Distributions for PPFPL
To further test the security of PPFPL in different data distributions, we evaluate the accuracy of PPFPL against feature attacks under different Non-IID conditions in MNIST, FMNIST and CIFAR10. In addition, we compare the performance of Krum, Foolsgold, ShieldFL, PBFL, FedDMC, and AntidoteFL. The specific experimental results are reported in Table I. Notably, since test average accuracy has fluctuated after each communication round, we select the average of five highest test average accuracies across the communication rounds. From Table I, we can observe that the performance of PPFPL is less affected by changes in data distribution, while other schemes are very susceptible to data distribution. This is attributed to that the client-submitted prototype does not change due to the change in data distribution. In addition, we observe a slight degradation in the performance of PPFPL with $Att=30\%$ attacks. This is because the tampered Non-IID data reduces the contribution of effective samples to the model, resulting in a slightly lower performance, which is a reasonable phenomenon. Therefore, our PPFPL improves performance in cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks, which is applicable to deployments in real-world scenarios.
TABLE I: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with feature attacks.
@@ VII-B3 Different Data Distributions for PPFPL
VII-B4 Secure Aggregation Protocol for PPFPL
To test the usefulness of aggregation protocol, we evaluate the performance of PPFPL with $\chi$ of -1, 0, 0.2, and 0.5, respectively, where $\chi=-1$ means that the protocol perform normalization verification and average aggregation, and the data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$). In addition, the $\boldsymbol{X}$ in Table II denotes that client-submitted prototypes are unnormalized and aggregated by simple averaging. As shown in Table II, we can observe from $\boldsymbol{X}$ that the performance of PPFPL suffers degradation to some extent compared to other settings. The performance degradation stems from the influence of malicious clients’ prototypes on the aggregation process, leading to a poisoned global prototype. However, the poisoned global prototype can only affect part of local model training, and cannot affect the minimization of local classification loss. Thus, in PPFPL, the model accuracy of benign clients is less susceptible to the influence of malicious clients. Furthermore, we observe that the presence of detection threshold improves the performance of PPFPL, which indicates that the secure aggregation protocol computes Byzantine-robust aggregation results, which reduces the influence of malicious clients and proves the effectiveness of the secure aggregation protocol. The above observation raises the question: whether PPFPL can maintain high performance in high proportion of attacks? We explore this question in the next experiment.
TABLE II: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with diffierent $\chi$ on feature attacks.
@@ VII-B4 Secure Aggregation Protocol for PPFPL
VII-B5 High Proportion Attacks for PPFPL
To test the performance of PPFPL under high proportion attacks, we evaluate it with Att = 20%, 40%, 60% and 80%, respectively. In addition, we set $\lambda$ as 0.01, 0.1 and 1 respectively, and the data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$). The $\lambda$ denotes the importance weight of the auxiliary term in loss function, which can be considered as the degree of influence among clients. The experimental results are shown in Fig. 7. We surprisingly observe that PPFPL still has high performance when Att is 60% or even 80%, which indicates that the training of benign clients is not interfered by malicious clients. This is because the benign client does not rely solely on information distributed by two servers in federated prototype learning, but relies heavily on its local training data. Furthermore, we observe that when Att is 20% or 40%, the accuracy increases slightly with the growth of $\lambda$ in the three datasets. On the contrary, when Att is 60% or 80%, the accuracy tends to decrease with the increase of $\lambda$. This is because the larger $\lambda$ strengthens the collaboration among clients, but makes PPFPL more vulnerable to data poisoning attacks from malicious clients. Conversely, the smaller value of $\lambda$ weakens collaboration among clients, but increases resistance to data poisoning attacks. Therefore, PPFPL can appropriately adjust the size of $\lambda$ according to actual conditions.
VII-B6 Stability of PPFPL

### v5

@@ 7 Experiments
7.1 Experimental Settings
7.1.1 Datasets and Models
Similar to previous works [16][38][55], we utilize three public available datasets, namely MNIST [56], FMNIST [57], and CIFAR10 [58], to evaluate performance of our PPFPL. Furthermore, we apply Convolution Neural Network (CNN) as local model to both MNIST and FMNIST. For CIFAR10, we employ ResNet18 as the local model, initialized with pre-trained parameters. These initial parameters have an accuracy of 27.5% on CIFAR10’s test set.
7.1.2 Hyper-parameters Settings of FL
We employ a cross-silo configuration in our FL experiments. Specifically, we set up 20 clients, each of which uploads local prototypes at each communication round. The number of rounds is set to 100, 150, and 150 for MNIST, FMNIST, and CIFAR10, respectively. The local learning rate is set to $\eta=0.01$, the importance weight to $\lambda=1$, the batch size to 64, and the number of local iterations to $E=5$. The detection threshold $\chi$ is configured as 0. Except for special declared, the above default settings are applied. These hyper-parameters are consistent across all clients. For encryption, we instantiate the CKKS in our experiments using TenSEAL [59] with the polynomial modulus degree of 8192 and a coefficient modulus of 200-bits, and a scale of 40-bits. This configuration offers 128-bits security level [60][61], and all encryption parameters follow the recommended settings provided by TenSEAL.
7.1.3 Non-IID Settings
To simulate the Non-IID data in cross-silo FL, we create class-space heterogeneity among clients, which is common in cross-silo scenarios. Specifically, large organizations (e.g., hospitals, companies) hold datasets with different classes, and their class distributions may differ significantly, or even be missing some classes altogether. When these organizations participate in federated training, the union of their data classes defines the entire FL classification task. This phenomenon leads to Non-IID data across organizations. However, the Dirichlet distribution [62] assumes that each client’s data is sampled from all classes, meaning each client usually has samples of every class. However, in cross-silo scenarios, some clients may have no samples from certain classes, which contradicts the assumptions of Dirichlet distribution.
To model the data distribution in cross-silo scenarios, similar to previous works [23][63], we define Avg as the mean number of data classes per client, and Std as the standard deviation of these class counts. In our experiments, we fix Avg to be 3 , 4 or 5, and fix Std to be 1 or 2, to create the class-space heterogeneity. Clients are randomly assigned to different classes, with partial class overlap among them. To visualize the different data distributions, we plot heat maps as shown in Fig. 2.
@@ 7.1.3 Non-IID Settings
Figure 2: Heat maps under different heterogeneity distributions.
7.1.4 Setting of Data Poisoning Attacks
@@
In our experiments, we evaluate two types of data poisoning attacks: feature attacks and label attacks. For feature attacks, malicious clients randomly alter their own training data features in a completely randomized manner without following any specific rule. For label attacks, malicious clients modify the labels of their training data to incorrect labels. All training data of malicious clients are modified. Unless otherwise specified, malicious clients keep their behaviors fixed during federated training, and the attack remains consistent throughout the training period. Since the number of malicious clients affects FL performance differently, we set different proportions of malicious clients, denotes as Att.
7.1.5 Evaluation Measure
@@
Our goal is to improve the performance of cross-silo PPFL under poisoned Non-IID data while maintaining robustness against data poisoning attacks. Therefore, we evaluate the FL performance by measuring the average accuracy of benign clients. For experiment comparison, we use FedAvg as the baseline, assuming it is not subject to data poisoning attacks. In addition, we compare PPFPL with robust schemes (i.e., Krum [37], Foolsgold [64], and FedDMC [38]) and privacy-preserving robust schemes (i.e., ShieldFL [19], PBFL [16], AntidoteFL [21], and FLDP-DM [65]) under the same experiment conditions.
7.2 Experimental Results
7.2.1 Visualization of Prototypes
@@
We visualize client-submitted prototypes in PPFPL using t-SNE, as shown in Fig. 3. Specifically, we display the distribution of prototypes from clients under feature and label attacks with $\textit{Att}=20\%$ on CIFAR10 within a single communication round. We observe that the prototypes of malicious clients deviate significantly from those of benign clients, regardless of features attacks or labels attacks. In addition, we notice that the prototypes submitted by malicious clients are concentrated in a ring area. This is because when malicious clients perform local model training, the features or labels are disrupted, causing the generated prototypes to spread from the center point to the surrounding areas. Most importantly, we observe a clear difference between the prototypes submitted by benign and malicious clients. This difference is prominent because it is independent of the data distribution. The finding validates the effectiveness of leveraging prototypes to defend against data poisoning attacks. To effectively filter out malicious prototypes, it is essential to employ a secure aggregation protocol that ensures Byzantine-robust aggregation results. The following is the security evaluation of PPFPL.
((a)) Feature attacks, Avg=3,Std=2.
((b)) Feature attacks, Avg=5,Std=1.
((c)) Label attacks, Avg=3,Std=2.
((d)) Label attacks, Avg=5,Std=1.
@@
Figure 3: The t-SNE visualization of client-submitted prototypes. Different colors indicate distinct classes. The symbol $\cdot$” denotes prototypes from benign clients, and $\times$” represents those from malicious clients.
7.2.2 Security Evaluation
@@
To evaluate the security of our framework, we evaluate PPFPL’s performance against feature attacks and label attacks on three datasets under Non-IID setting with Avg=3 and Std=2. Specifically, the proportion of malicious clients is set to Att=20%. Furthermore, we compare PPFPL with existing defense schemes (i.e., FedAvg, ShieldFL, PBFL, FedDMC, AntidoteFL, and FLDP-DM), where the global model learned by FedAvg is not subject to data poisoning attacks.
The experiment results are shown in Fig. 4, where the black dashed line represents the highest accuracy of PPFPL without attacks. We observe that ShieldFL and PBFL suffer significant accuracy degradation under attacks compared to FedAvg without attacks on the CIFAR10 dataset. This indicates that as datasets become more complex, the negative impact of data poisoning attacks on model accuracy becomes more obvious. This is mainly because complex datasets increase the difficulty of defense, and poisoned Non-IID data further degrades the defense performance. Notably, FedDMC and AntidoteFL achieve higher performance compared to FedAvg because they adjust the aggregation weights of model updates, which mitigates the bias caused by Non-IID data and improves the performance of federated learning. Moreover, PPFPL (under attacks) outperforms FedAvg (no attacks), with its accuracy approaching the highest accuracy of PPFPL without attacks when it nears convergence. This demonstrates that PPFPL guarantees the high accuracy of the learned model in the presence of poisoned Non-IID data, which is benefited that the prototype is not affected by tampered data distribution, while the secure aggregation protocol resists malicious prototypes submitted by clients. Thus, our PPFPL overcomes the difficulty confronted by these defense schemes, and improves the FL performance in poisoned Non-IID data.
Additionally, we evaluate the change of loss during training for PPFPL and the compared schemes, as shown in Fig. 5. We observe that PBFL’s loss fluctuates significantly throughout the training process, while PPFPL demonstrates relatively smooth convergence. This indicates that data poisoning attacks severely disrupt PBFL’s convergence, while the convergence of PPFPL is not affected by data poisoning attacks, thereby preserving model accuracy. This is consistent with our Corollary 3. Specifically, PPFPL still converges in the presence of data poisoning attacks as long as $\lambda$, $E$, and $\eta$ satisfy a specific relationship among them. Moreover, PPFPL’s loss at convergence is lower than other defense schemes under CIFAR10, which can satisfy the design goal of security. Notably, our experiments show that feature attacks and label attacks have similar impacts on the performance of PPFPL and other schemes. This indicates that the impact on model performance is similar under poisoned Non-IID data, regardless of the type of data poisoning attacks.
@@
7.2.3 Different Data Distributions for PPFPL
To further test the security of PPFPL under different data distributions, we evaluate the accuracy of PPFPL against feature attacks under different Non-IID conditions in MNIST, FMNIST and CIFAR10. In addition, we compare its performance with Krum, Foolsgold, ShieldFL, PBFL, FedDMC, AntidoteFL, and FLDP-DM. The specific experimental results are reported in TABLE 2. Notably, since test average accuracy has fluctuated after each communication round, we select the average of five highest test average accuracies across all communication rounds. From TABLE 2, we can observe that the performance of PPFPL is not easily affected by changes in data distribution, while other schemes are very susceptible to data distribution. This is attributed to that the client-submitted prototype does not change due to the change in data distribution. In addition, we observe a slight degradation in the performance of PPFPL with $Att=30\%$ attacks. This is because the tampered Non-IID data reduces the contribution of effective samples to the model, resulting in a slightly lower performance, which is a reasonable phenomenon. Overall, our PPFPL improves performance in cross-silo PPFL under poisoned Non-IID data while resisting data poisoning attacks, making it suitable for real-world deployments.
Table 2: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with feature attacks.
@@ 7.2.3 Different Data Distributions for PPFPL
7.2.4 Secure Aggregation Protocol for PPFPL
To test the effectiveness of the aggregation protocol, we evaluate the performance of PPFPL with $\chi$ of -1, 0, 0.2, and 0.5, respectively. Here, $\chi=-1$ means that the protocol performs normalization verification and average aggregation. The data distribution is set to Non-IID (i.e., Avg=3, Std=2). In addition, $\boldsymbol{X}$ in Table 3 denotes that client-submitted prototypes are unnormalized and aggregated using simple averaging. As shown in TABLE 3, we can observe from $\boldsymbol{X}$ that the performance of PPFPL suffers degradation to some extent compared to other settings. The performance degradation is because malicious prototype poisons the global prototype. However, the poisoned global prototype can only affect part of local model training, and cannot affect the minimization of local classification loss. Thus, the model accuracy among benign clients in PPFPL remains relatively robust against the influence of malicious clients. Furthermore, we observe that the presence of detection threshold improves the performance of PPFPL, which indicates that the secure aggregation protocol computes Byzantine-robust aggregation results. The above observation raises the question: whether PPFPL can maintain high performance in high proportion of attacks? We address this question in the subsequent experiments.
Table 3: Test Average Accuracy (%) on MNIST, FMNIST and CIFAR10 with diffierent $\chi$ on feature attacks.
@@ 7.2.4 Secure Aggregation Protocol for PPFPL
7.2.5 High Proportion Attacks for PPFPL
To evaluate the performance of PPFPL under high proportion attacks, we test it with Att = 20%, 40%, 60% and 80%. In addition, we set $\lambda$ as 0.01, 0.1 and 1 under Non-IID data distribution (Avg=3, Std=2). The $\lambda$ denotes the importance weight of the auxiliary term in the loss function, which can be considered as the degree of influence among clients. The experimental results are shown in Fig. 6. Surprisingly, We observe that PPFPL still has high performance when Att is 60% or even 80%, which indicates that the training of benign clients is not interfered by malicious clients. This is because the benign clients do not rely solely on information distributed by two servers in federated prototype learning, but relies heavily on its local training data. Furthermore, we observe that when Att is 20% or 40%, the accuracy increases slightly with the growth of $\lambda$ in the three datasets. On the contrary, when Att is 60% or 80%, the accuracy tends to decrease with the increase of $\lambda$. This is because the larger $\lambda$ strengthens the collaboration among clients, but makes PPFPL more vulnerable to data poisoning attacks from malicious clients. Conversely, the smaller value of $\lambda$ weakens collaboration among clients, but enhances resistance to data poisoning attacks. Therefore, PPFPL can appropriately adjust the size of $\lambda$ according to actual conditions.
7.2.6 Dynamic Data Poisoning Attacks for PPFPL
In practical cross-silo scenarios, the behavior of malicious clients may vary across different rounds. To emulate this scenario, we configure each malicious client to alternate between feature attacks and label attacks in consecutive rounds, termed dynamic poisoning attacks. We evaluate the performance of PPFPL under the dynamic attacks. The experimental results (shown in TABLE 4) demonstrate that the performance of PPFPL only slightly decreases with the increase in the number of malicious clients, mainly due to the reduction in benign samples. Thus, PPFPL maintains superior performance under dynamic poisoning attacks. This is because, regardless of how malicious clients manipulate their submitted prototypes, our secure aggregation protocol can limit their impact to the directional perturbation. The protocol filters out prototypes with excessive directional deviation, ensuring that PPFPL is resilient to dynamic poisoning attacks.
Table 4: Test Average Accuracy (%) of PPFPL against dynamic poisoning attacks.
Datasets
Data
Distributions
Proportion of Malicious Clients (%)
0% 10% 20% 30% 40%
CIFAR10 Avg=3, Avg=1 83.05 83.20 82.79 82.27 81.55
Avg=3, Avg=2 83.29 83.13 82.83 82.04 81.26
Avg=4, Avg=1 82.04 81.87 80.71 80.93 80.60
Avg=4, Avg=2 82.11 82.06 82.45 81.75 81.21
Avg=5, Avg=1 77.82 77.31 77.24 76.92 76.49
Avg=5, Avg=2 77.95 78.12 78.03 77.80 76.22
7.2.7 Stability of PPFPL